# Solving Inverse Problems in Protein Space Using Diffusion-Based Priors

### v1

@@ 4.1 General Approach
While most of previous works leveraging a diffusion model for inverse problems aim at sampling from the posterior distribution $p(\mathbf{x}|\mathbf{Y})$, we are interested here in scenarios where the measurements convey enough information to make the MAP estimate unique and well-defined. In the results section, we show that MAP estimation outperforms posterior sampling in such well-defined problems.
@@
$\hat{\mathbf{z}}_{0}$ $=\operatorname*{argmin}_{\mathbf{z}}\frac{\rho}{2}\|\mathbf{z}-\tilde{\mathbf{z}}_{0}\|_{2}^{2}-\sum_{i=1}^{n}\log p_{0}(\mathbf{y}_{i}|\mathbf{z})$ Maximize likelihood,
@@
Here, no specific assumptions have been made on the likelihood term and this framework could hypothetically be applied on any set of measurements for which we have a physics-based model of the measurement process. Since the second step is not tractable in most cases, we replace the explicit minimization with a gradient step with momentum from the iterate $\tilde{\mathbf{z}}_{0}$. This step can be implemented efficiently using automatic differentiation.
@@
Diffusion model: correlation matrix $\mathbf{R}$, denoiser $\hat{\mathbf{x}}_{\theta}(\mathbf{x},t)$, schedule $\{\alpha_{t}\}_{t=1}^{T}$.
@@
     $\forall i,\mathbf{v}_{i}=\rho_{i}\mathbf{v}_{i}+\lambda_{i}\nabla_{\mathbf{z}}f_{i}(\mathbf{y}_{i},\mathbf{z})|_{\mathbf{z}=\tilde{\mathbf{z}}_{0}}$ Accumulate gradient of log likelihood
@@ 4.2 Preconditioning for Linear Measurements
Figure 2: Condition number of $\mathbf{R}$ vs. number of residues $N$.
and $C$ does not depend on $\mathbf{z}$. As shown in Figure 2, the condition number of $\mathbf{R}$ (i.e., the ratio between its largest and smallest singular values) grows as a power function of the number of residues. For typical proteins ($N\geq 100$), this condition number exceeds $100$, making the maximization of the above term an ill-conditioned problem. In order to make gradient-based optimization more efficient, we propose to precondition the problem by precomputing a singular value decomposition $\mathbf{A}\mathbf{R}=\mathbf{U}\mathbf{S}\mathbf{V}^{T}$ and to set $\mathbf{\Sigma}=\sigma^{2}\mathbf{U}\mathbf{S}\mathbf{S}^{T}\mathbf{U}^{T}$. Note that this is equivalent to modeling the measurement process as $\mathbf{y}=\mathbf{A}\mathbf{R}(\mathbf{z}+\tilde{\eta})$ with $\tilde{\eta}\sim\mathcal{N}(0,\sigma^{2})$. In other words, we assume that the noise $\eta$ preserves the simple patterns in proteins, which is a reasonable hypothesis if, for example, $\mathbf{y}$ is an incomplete atomic model obtained by an upstream reconstruction algorithm that leverages prior knowledge on protein structures. The log-likelihood then becomes
@@ 4.3 Application to Atomic Model Building
where $\Gamma$ is an operator that places a sum of 5 isotropic Gaussians centered on each heavy atom. The amplitudes and standard deviations of these Gaussians, known as “form factors”, are tabulated [24] and depend on the chemical element they are centered on. $\mathcal{B}$ represents the effect of “B-factors” [34] and can be viewed as a spatially-dependent blurring kernel modelling molecular motions and/or signal damping by the transfer function of the electron microscope. $\eta$ models isotropic Gaussian noise of variance $\sigma^{2}$. This measurement model leads to the following log-likelihood:
@@
$\log p_{0}(\mathbf{y},\mathbf{s},\bar{\mathbf{x}}|\mathbf{x})=\log p_{0}(\mathbf{s}|\mathbf{x})+\log p_{0}(\bar{\mathbf{x}}|\mathbf{x})+\log p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s}).$ (13)
On the right-hand side, the first term can be approximated using the learned conditional distribution $p_{\theta}(\mathbf{s}|\mathbf{x},t=0)$. We model $\bar{\mathbf{x}}=\mathbf{M}\mathbf{x}+\eta$ so that the second term can be handled by the preconditioning procedure described in the previous section. Finally, the last term involves the marginalization of $p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s},\mathbf{\chi})$ over $\chi$. This marginalization is not tractable but (12) provides a lower bound:
$\log p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s})\geq\mathbb{E}_{\mathbf{\chi}\sim p_{0}(\mathbf{\chi}|\mathbf{x},\mathbf{s})}\Bigl{[}\log p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s},\mathbf{\chi})\Bigr{]}\approx\mathbb{E}_{\mathbf{\chi}\sim p_{\theta}(\mathbf{\chi}|\mathbf{x},\mathbf{s},t=0)}\Bigl{[}\log p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s},\mathbf{\chi})\Bigr{]},$ (14)
using Jensen’s inequality. The expectation is approximated by Monte Carlo sampling and gradients of $\chi$ with respect to $\mathbf{x}$ are computed by automatic differentiation through the autoregressive sampler of $\mathbf{\chi}$.
@@ 5 Experiments
Our method uses the publicly released version of Chroma [^1] https://github.com/generatebio/chroma [29]. We run all our experiments using structures of proteins downloaded from the Protein Data Bank (PDB) [7]. In order to select proteins that do not belong to the training dataset of Chroma, here we only consider structures that were released after 2022-03-20 (Chroma was trained on a filtered version of the PDB queried on that date). We only consider single-chain structures for which all the residues have been spatially resolved, so that the Root Mean Square Deviation (RMSD) of the predicted Cartesian coordinates can be computed for all the heavy atoms in the backbone. In each experiment, we run 8 replicas in parallel on a single NVIDIA A100 GPU.
@@
Given an incomplete atomic model of a protein, our first task is to predict the coordinates of all heavy atoms in the backbone. The forward measurement process can be modeled as $\mathbf{y}=\mathbf{M}\mathbf{x}$ where $\mathbf{M}\in\{0,1\}^{(4N/k)\times 4N}$ is a masking matrix ($\mathbf{M}\mathbf{1}=\mathbf{1}$) and $k$ is the subsampling factor. We consider the case where, for each residue, the location of all 4 heavy atoms on the backbone (N, Cα, C, O) is either known or unknown. Residues of known locations are regularly spaced along the backbone every $k$ residues. We compare our results to the baseline Chroma conditioned with a SubstructureConditioner [29]. This baseline samples from the posterior probability $p(\mathbf{x}|\mathbf{y})$ using unadjusted Langevin dynamics. We use 1000 diffusion steps for our method and the baseline.
In Figure 3, we show our results on ATAD2 (PDB:7qum[15, 5], a cancer-associated protein of 130 residues. The protein was resolved at a resolution of 1.5 Å using X-ray crystallography. Our method recovers the target structure without loss of information ($\text{RMSD}<1.5~{}\text{\AA}$) for subsampling factors of 2, 4 and 8. Fig. 3.b shows that our method outperforms the baseline and highlights the importance of the diffusion-based prior. When the subsampling factor is large ($\geq 32$), the reconstruction accuracy decreases but the method inpaints unknown regions with realistic secondary structures (see quantitative evaluation in the supplementary). Note that making the conditioning information sparser (increasing the subsampling factor) tends to close the gap between our method (MAP estimation) and the baseline (posterior sampling). We provide results on four other protein structures in the supplementary.
Figure 4: Atomic Model Refinement. Results on the TecA bacterial toxin (PDB:7pzt, 160 residues). (a) Qualitative results. From left to right: the input density map at 2.0 Å resolution, the incomplete model given by ModelAngelo and our refined models (1 sample and 5 samples), overlaid on the target structure in transparency. (b) RMSD of alpha carbons vs. completeness (number of predicted residues / total number of residues) with ModelAngelo (MA) and our method. We run 5 experiments and report the mean of the lowest RMSD on $\alpha$-carbons over 8 replicas (±$\pm 1$ std). The experimental (deposited) resolution is indicated with a dashed line.
@@
Next, we evaluate our method on the model refinement task. We use ChimeraX [48, 62] to simulate the 3D density maps of a single-chain protein, at different resolutions. We run the state-of-the-art model building method ModelAngelo [31] on the simulated maps, using the known sequence. Then, we provide ModelAngelo’s output (an incomplete model) to our method, along with the density map and the sequence. We use ModelAngelo’s default parameters. To evaluate our method, we report the RMSD of the predicted structure for the $X$most well-resolved alpha carbons (compared to the deposited structure), for $X\in[0,100]$ ($X$ is called the “completeness”).
In Figure 4, we show our results on a synthetic density map of the TecA bacterial toxin (PDB:7pzt). This protein structure belongs to the CASP15 dataset, released in May 2022 to evaluate protein structure modeling methods. Proteins belonging to this dataset were selected for not having close homologous in the PDB, which ensures that Chroma was not trained on similar structures. The protein was resolved at 1.84 Å resolution using X-ray crystallography. Note that we do not chose a structure that was resolved with cryo-EM because most cryo-EM-resolved models are incomplete. For both input resolutions (2.0 Å and 4.0 Å), our method improves on ModelAngelo’s accuracy for a fixed completeness level (and reaches a higher completeness for the same accuracy). We explore the effect of removing some of the conditioning information in the supplementary materials.
@@
Finally, we assume we are given a set of pairwise distances between alpha carbons of a protein and we use our method to predict a full 3D structure. This task is a simplification of the reconstruction problem in paramagnetic NMR spectroscopy, where one can obtain information about the relative distances and orientations between pairs of atoms via the nuclear Overhauser effect and sparse paramagnetic restraints, and must deduce the Cartesian coordinates of every atom [38, 39]. Formally, our measurement model is $\mathbf{y}=\|\mathbf{D}\mathbf{x}\|_{2}\in\mathbb{R}^{m}$, where $\mathbf{D}\in\{-1,0,1\}^{m\times 4N}$ is the distance matrix and the norm is taken row-wise (in $xyz$ space). $\mathbf{D}$ contains a single “1” and a single “-1” in each row and is not redundant (the distance between a given pair of atoms is measured at most once). $m$ corresponds to the number of measured distances.
We evaluate our method on the bromodomain-containing protein 4 (BRD4, PDB:7r5b [69]), a protein involved in the development of a specific type of cancer (NUT midline carinoma) [22] and targeted by pharmaceutical drugs [14]. For a given number $m$, we randomly sample $m$ pairs of alpha carbons (without redundancy) between which we assume the distances to be known. Our results are shown in Figure 5. When 500 pairwise distances or more are known, our method recovers the structural information of the target structure ($\text{RMSD}<1.77~{}\text{\AA}$, the resolution of the deposited structure resolved with X-ray crystallography). We conduct the same experiment without the diffusion model and show a drop of accuracy, highlighting the importance of the generative prior. Note that, when the diffusion model is removed, increasing the number of measurements increases the number of local minimima in the objective function and can therefore hurt the reconstruction quality (plot in Fig. 5, orange curve, far-right part).
@@ 6 Discussion
We focused here on cases where the measurement process is simulated and perfectly known. Considering real experimental measurements (e.g., cryo-EM or NMR data) would raise complex and exciting challenges. In particular, we hope ADP-3D can serve as a precursor for reconstructing all-atom models directly from cryo-EM images, a key to turn empirical conformational distributions into thermodynamic information [21]. In cases where the measurement process cannot be faithfully modeled due to complex nonidealities, or when the measurement process is not differentiable, our framework reaches its boundaries. At the detriment of versatility, exploring the possibility of finetuning a pretrained diffusion model on paired data for conditional generation constitutes another promising avenue for future work.
Acknowledgements.
The author(s) are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is a consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Office of Information Technology’s Research Computing. This material is based upon work supported by the National Science Foundation under award number 2303178 to SFK. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
We thank Mike Dunne and Jay Shenoy for their insightful feedback on this manuscript.
@@ Appendix A Log-Likelihood Functions
$\Gamma$ is an operator turning an all-atom model into a density map. Given $\mathbf{X}=(\mathbf{x},\mathbf{s},\mathbf{\chi})=[X_{1},\ldots,X_{A}]^{T}\in(\mathbb{R}^{3})^{A}$, the Cartesian coordinates of $A$ heavy atoms,
$\Gamma(\mathbf{X}):x\in\mathbb{R}^{3}\mapsto\sum_{i=1}^{A}\sum_{j=1}^{5}4\pi\frac{a_{i,j}}{b_{i,j}}\exp\left(-\frac{4\pi^{2}}{b_{i,j}}\|X_{i}-x\|_{2}^{2}\right)\in\mathbb{R},$ (17)
where $a_{i,j}$ and $b_{i,j}$ are tabulated form factors [24]. The norm $\|\mathbf{y}-\Gamma(\mathbf{X})\|_{2}^{2}$ is computed directly in Fourier space using Parseval’s theorem, up to a time-dependent resolution $r_{t}$. Here, we assume the B-factors are unknown and we neglect their contribution in the log-likelihood function as we consider high-resolution density maps ($\leq 4~{}\text{\AA}$).

### v2

@@ 4.1 General Approach
While most of previous works leveraging a diffusion model for inverse problems aim at sampling from the posterior distribution $p(\mathbf{x}|\mathbf{Y})$, we are interested here in scenarios where the measurements convey enough information to make the MAP estimate unique and well-defined.
@@
$\hat{\mathbf{z}}_{0}$ $=\operatorname*{argmin}_{\mathbf{z}}\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}}{2}\|\mathbf{z}-\tilde{\mathbf{z}}_{0}\|_{2}^{2}-\sum_{i=1}^{n}\log p_{0}(\mathbf{y}_{i}|\mathbf{z})$ Maximize likelihood,
@@
Here, no specific assumptions have been made on the likelihood term and this framework could hypothetically be applied on any set of measurements for which we have a physics-based model of the measurement process. Since the second step is not tractable in most cases, we replace the explicit minimization with a gradient step with momentum from the iterate $\tilde{\mathbf{z}}_{0}$. This step can be implemented efficiently using automatic differentiation. The gradient of $\|\mathbf{z}-\hat{\mathbf{z}}_{0}\|_{2}^{2}$ w.r.t $\mathbf{z}$ in $\hat{\mathbf{z}}_{0}$ being null, the method does not depend on the choice $\gamma$.
@@
Diffusion model: correlation matrix $\mathbf{R}$, denoiser $\hat{\mathbf{x}}_{\theta}(\mathbf{x},t)$, schedule $\{\alpha_{t},\sigma_{t}\}_{t=1}^{T}$.
@@
     $\forall i,\mathbf{v}_{i}=\rho_{i}\mathbf{v}_{i}+\lambda_{i}\nabla_{\mathbf{z}}f_{i}(\mathbf{y}_{i},\mathbf{z})|_{\mathbf{z}=\tilde{\mathbf{z}}_{0}}$ Accumulate gradient of log-likelihood
@@ 4.2 Preconditioning for Linear Measurements
and $C$ does not depend on $\mathbf{z}$. As shown in the supplements, the condition number of $\mathbf{R}$ (i.e., the ratio between its largest and smallest singular values) grows as a power function of the number of residues. For typical proteins ($N\geq 100$), this condition number exceeds $100$, making the maximization of the above term an ill-conditioned problem. In order to make gradient-based optimization more efficient, we propose to precondition the problem by precomputing a singular value decomposition $\mathbf{A}\mathbf{R}=\mathbf{U}\mathbf{S}\mathbf{V}^{T}$ and to set $\mathbf{\Sigma}=\sigma^{2}\mathbf{U}\mathbf{S}\mathbf{S}^{T}\mathbf{U}^{T}$. Note that this is equivalent to modeling the measurement process as $\mathbf{y}=\mathbf{A}\mathbf{R}(\mathbf{z}+\tilde{\eta})$ with $\tilde{\eta}\sim\mathcal{N}(0,\sigma^{2})$. In other words, we assume that the noise $\eta$ preserves the simple patterns in proteins, which is a reasonable hypothesis if, for example, $\mathbf{y}$ is an incomplete atomic model obtained by an upstream reconstruction algorithm that leverages prior knowledge on protein structures. The log-likelihood then becomes
@@ 4.3 Application to Atomic Model Building
where $\Gamma$ is an operator that places a sum of 5 isotropic Gaussians centered on each heavy atom. The amplitudes and standard deviations of these Gaussians, known as “form factors”, are tabulated [32] and depend on the chemical element they are centered on. $\mathcal{B}$ represents the effect of “B-factors” [40] and can be viewed as a spatially dependent blurring kernel modelling molecular motions and/or signal damping by the transfer function of the electron microscope. $\eta$ models isotropic Gaussian noise of variance $\sigma^{2}$. This measurement model leads to the following log-likelihood:
@@
$\log p_{0}(\mathbf{y},\mathbf{s},\bar{\mathbf{x}}|\mathbf{x}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\log p_{0}(\bar{\mathbf{x}}|\mathbf{x})+p_{0}(\mathbf{y},\mathbf{s}|\mathbf{x})}=\log p_{0}(\bar{\mathbf{x}}|\mathbf{x})+\log p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s})+\log p_{0}(\mathbf{s}|\mathbf{x}).$ (13)
On the right-hand side, the last term can be approximated using the learned conditional distribution $p_{\theta}(\mathbf{s}|\mathbf{x})$. We model $\bar{\mathbf{x}}=\mathbf{M}\mathbf{x}+\eta$ so that the first term can be handled by the preconditioning procedure described in the previous section. Finally, the middle term involves the marginalization of $p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s},\mathbf{\chi})$ over $\chi$. This marginalization is not tractable but (12) provides a lower bound:
$\log p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s})\geq\mathbb{E}_{\mathbf{\chi}\sim p_{0}(\mathbf{\chi}|\mathbf{x},\mathbf{s})}\Bigl{[}\log p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s},\mathbf{\chi})\Bigr{]}\approx\mathbb{E}_{\mathbf{\chi}\sim p_{\theta}(\mathbf{\chi}|\mathbf{x},\mathbf{s})}\Bigl{[}\log p_{0}(\mathbf{y}|\mathbf{x},\mathbf{s},\mathbf{\chi})\Bigr{]},$ (14)
using Jensen’s inequality. The expectation is approximated by Monte Carlo sampling and gradients of $\chi$ with respect to $\mathbf{x}$ are computed by automatic differentiation through the autoregressive sampler of $\mathbf{\chi}$, following the “reparameterization trick” [43].
@@ 5 Experiments
Our main results are obtained using the publicly released version of Chroma [^1] https://github.com/generatebio/chroma [37]. We provide additional results with the publicly released version of RFdiffusion [^2] https://github.com/RosettaCommons/RFdiffusion [79] in the supplements. We run all our experiments using structures of proteins downloaded from the Protein Data Bank (PDB) [6]. In order to select proteins that do not belong to the training dataset of Chroma, here we only consider structures that were released after 2022-03-20 (Chroma was trained on a filtered version of the PDB queried on that date). We provide additional results on structures taken from the CASP15 dataset in the supplements. For the proteins that are not fully modeled on the PDB, we mask out the residues with ground truth coordinates before computing the Root Mean Square Deviation (RMSD). In each experiment, we run 8 replicas in parallel on a single NVIDIA A100 GPU. Further details about each target structure are provided in the supplements.
@@
Given an incomplete atomic model of a protein, our first task is to predict the coordinates of all heavy atoms in the backbone. This first task is designed as a toy problem, with no immediate application to real data, to validate and evaluate our method. The forward measurement process can be modeled as $\mathbf{y}=\mathbf{M}\mathbf{x}$ where $\mathbf{M}\in\{0,1\}^{(4N/k)\times 4N}$ is a masking matrix ($\mathbf{M}\mathbf{1}=\mathbf{1}$) and $k$ is the subsampling factor. We consider the case where, for each residue, the location of all 4 heavy atoms on the backbone (N, Cα, C, O) is either known or unknown. Residues of known locations are regularly spaced along the backbone every $k$ residues. We compare our results to the baseline Chroma conditioned with a SubstructureConditioner [37]. This baseline samples from the posterior probability $p(\mathbf{x}|\mathbf{y})$ using unadjusted Langevin dynamics. We use 1000 diffusion steps for our method and the baseline.
In Figure 2, we show our results on ATAD2 (PDB:7qum[21, 4], a cancer-associated protein of 130 residues. The protein was resolved at a resolution of 1.5 Å using X-ray crystallography. Our method recovers the target structure without loss of information ($\text{RMSD}<1.5~{}\text{\AA}$) for subsampling factors of 2, 4 and 8. Fig. 2.b shows that our method outperforms the baseline and highlights the importance of the diffusion-based prior. When the subsampling factor is large ($\geq 32$), the reconstruction accuracy decreases but the method inpaints unknown regions with realistic secondary structures (see quantitative evaluation in the supplementary). Note that making the conditioning information sparser (increasing the subsampling factor) tends to close the gap between our method (MAP estimation) and the baseline (posterior sampling).
Figure 3: Atomic Model Refinement. (a) Experimental density map, ModelAngelo’s incomplete model and ADP-3D’s atomic model. (b) RMSD of alpha carbons vs. completeness (number of predicted residues / total number of residues) with ModelAngelo (MA) and our method. For EMD-26482 (third row), we remove uniformly sampled residues from the ModelAngelo model until reaching sub-80% of completeness. The RMSD is computed with respect to the deposited structure on the PDB.
@@
[…]
We show qualitative and quantitative results in Figure 3. For all structures, ADP-3D improves the accuracy of the ModelAngelo model for the same level of completeness. Note that the RMSD can only be computed up to the completeness level of the deposited structure. We provide additional results in the supplements and investigate on the influence of the resolution of the input density map and we perform an ablation study on the conditioning information.
@@
Finally, we assume we are given a set of pairwise distances between alpha carbons and we use our method to predict a full 3D structure. This task is a simplification of the reconstruction problem in paramagnetic NMR spectroscopy, where one can obtain information about the relative distances and orientations between pairs of atoms via the nuclear Overhauser effect and sparse paramagnetic restraints, and must deduce the Cartesian coordinates of every atom [44, 45], [64, 80, 54]. Formally, our measurement model is $\mathbf{y}=\|\mathbf{D}\mathbf{x}\|_{2}\in\mathbb{R}^{m}$, where $\mathbf{D}\in\{-1,0,1\}^{m\times 4N}$ is the distance matrix and the norm is taken row-wise (in $xyz$ space). $\mathbf{D}$ contains a single “1” and a single “-1” in each row and is not redundant (the distance between a given pair of atoms is measured at most once). $m$ corresponds to the number of measured distances.
We evaluate our method on the bromodomain-containing protein 4 (BRD4, PDB:7r5b [78]), a protein involved in the development of a specific type of cancer (NUT midline carinoma) [28] and targeted by pharmaceutical drugs [18]. For a given number $m$, we randomly sample $m$ pairs of alpha carbons (without redundancy) between which we assume the distances to be known. Our results are shown in Figure 4. When 500 pairwise distances or more are known, our method recovers the structural information of the target structure ($\text{RMSD}<1.77~{}\text{\AA}$, the resolution of the deposited structure resolved with X-ray crystallography). We conduct the same experiment without the diffusion model and show a drop of accuracy, highlighting the importance of the generative prior. Note that, when the diffusion model is removed, increasing the number of measurements increases the number of local minima in the objective function and can therefore hurt the reconstruction quality (plot in Fig. 4, orange curve, far-right part).
@@ 6 Discussion
Considering real data (e.g., cryo-EM, X-ray crystallography or NMR data) raises complex and exciting challenges, as experimental measurements of any one specific task typically require a lot of domain-specific processing. For example, in a real scenario, NMR experiments cannot probe long pairwise distances above 6 Å. Taking these constraints into account and applying ADP-3D to real NMR data would be an exciting direction for future work. In cryo-EM, most of the analyzed proteins are multi-chain while the current implementation of ADP-3D only supports single-chain structures. Extending our framework to multi-chain structures, which would possible using diffusion models like Chroma, would be an impactful future direction.
In cases where the measurement process cannot be faithfully modeled due to complex nonidealities, or when the measurement process is not differentiable, our framework reaches its boundaries. Exploring the possibility of finetuning a pretrained diffusion model on paired data for conditional generation constitutes another promising avenue for future work.
Acknowledgment
The author(s) are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is a consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Office of Information Technology’s Research Computing. The Zhong lab gratefully acknowledges support from the Princeton Catalysis Initiative, Princeton School of Engineering and Applied Sciences, and Generate Biomedicines. The funders had no role in study design, data collection and analysis, decision to publish or preparation of the manuscript.
Code Availability
Our code is publicly available at: https://github.com/axel-levy/axlevy-adp-3d
@@ Appendix A Log-Likelihood Functions
$\Gamma$ is an operator turning an all-atom model into a density map with a user-defined resolution, following the pdb2mrc implementation of EMAN [50]. Given $\mathbf{X}=(\mathbf{x},\mathbf{s},\mathbf{\chi})=[X_{1},\ldots,X_{A}]^{T}\in(\mathbb{R}^{3})^{A}$, the Cartesian coordinates of $A$ heavy atoms,
$\Gamma(\mathbf{X}):x\in\mathbb{R}^{3}\mapsto\sum_{i=1}^{A}Z_{i}\exp\left(-\frac{\|X_{i}-x\|_{2}^{2}}{2s^{2}}\right)\in\mathbb{R},$ (17)
where $Z_{i}$ is the atomic number of atom $i$ and $s=\text{resolution}/(\sqrt{2}\pi)$. The norm $\|\mathbf{y}-\Gamma(\mathbf{X})\|_{2}^{2}$ is computed directly in Fourier space using Parseval’s theorem, up to a time-dependent resolution $r_{t}$.